OpenStudy (anonymous):

if you had 12 1/2 yrds of material it takes 3 5/6 for 1 cape .how many capes can i make

OpenStudy (anonymous):

5

OpenStudy (anonymous):

[12.5/(23/6)] = 3.25 Therefore you can only make 3 capes.

OpenStudy (anonymous):

How did you get 5?

OpenStudy (anonymous):

3(5/6)=2.5 2.5*5=12

OpenStudy (anonymous):

I thought 3 5/6 was a mixed fraction. That is not clear.

OpenStudy (anonymous):

Oh you are mostprobably correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\frac{12+\frac{1}{2}}{3+\frac{5}{6}}=\frac{\frac{25}{2}}{\frac{23}{6}}=\frac{25}{2}*\frac{6}{23}=\frac{150}{46} \]\[\text{Floor}\left[\frac{150}{46}\right]=3\text{ capes} \]